Martine leaves home at 9 a.m., bicycling at a rate of 24 mi/hr. Two hours later, John leaves, driving at the rate of 48 mi/hr. At what time will John catch up with Martine?
Answer by checkley77(12844)

D=RT
D=24T FOR MARTINE
D=48(T-2) FOR JOHN
24T=48(T-2)
24T=48T-96
24T-48T=-96
-24T=-96
T=-96/-24
T=4 HOURS AFTER 9:00 AM THEY WILL MEET.